# High-dimensional dynamics in low-dimensional networks

### v4

@@ Appendix A Comparison to previous work on dynamics of low-rank networks
The proof of Claim 2 also tells us the dominant directions of variability whenever dynamics are low-dimensional as in Figure 2 and Figure 3g. Since $\bf{z}$} is approximately aligned to the $U$ in the proof, we may conclude that low-dimensional dynamics are caused by excess variability along the column space of $U$, i.e., the column space of $W_{0}$.
The conclusions above were reached by assuming that ${\bf{x}}(t)$ varies more slowly than $\tau$. However, note that Eq. (13) shows that the same conclusions can be reached in the context of variability at any frequency mode, $f$, regardless of how quickly ${\bf{x}}(t)$ varies. Specifically, variability in ${\bf{z}}(t)$ at a particular frequency mode (as quantified by $C^{{\bf{z}}}(f)$) is defined in terms of a regularized inverse of $W$, just like the stationary variance in Eq. (15). The only salient difference is that the regularizer is a scalar multiple of the identity, $\widetilde{r}(f)I$, in Eq. (13) instead of the identity itself, as in Eq. (15). This difference could be important, for example, if $\widetilde{r}(f)=0$ at some frequency, $f$, for which $C^{{\bf{x}}}(f)$ is not close to zero (corresponding to a situation in which variability in ${\bf{x}}(t)$ is faster than $\tau$). Next, in Appendix C, we generalize away from the assumption of slow inputs using a Laplace transform.
Appendix C The analysis of fast and transient dynamics.
Above, we considered slow or static perturbations, ${\bf{x}}(t)$. Specifically, we assumed that perturbations were constant in time or changed more slowly than the intrinsic dynamics of the network (quantified by $\tau$ in our models). This assumption applies to all of the examples used in the text, but in some applications, perturbations are faster than intrinsic network dynamics. We now consider transient perturbations, ${\bf{x}}(t)$, with arbitrary timescales. We can no longer rely on the quasi-steady state approximation from Eq. (2) in this case. Instead, we take the Laplace transform in Eq. (6) to obtain
$\hat{{\bf{z}}}=\hat{H}\hat{{\bf{z}}}+\hat{B}\hat{{\bf{x}}}$ (18)
where $\hat{\psi}(s)=\int\psi(t)e^{-st}dt$ is the Laplace transform of $\psi(t)$ and
$\hat{H}(s)$ $=(1+\tau s)^{-1}W$ (19)
$\hat{B}(s)$ $=(1+\tau s)^{-1}.$
Solving the implicit Eq. (18) for $\hat{{\bf{z}}}$ gives the response to arbitrary inputs in the Laplace domain,
$\hat{{\bf{z}}}=[I-\hat{H}]^{-1}\hat{B}\hat{{\bf{x}}}$. (20)
The $N\times N$ matrix $\hat{H}(s)$ can be interpreted as a measure of the effective connectivity of the linearized network at mode $s$ because each entry $\hat{H}_{jk}(s)$, represents the linearized response of ${\bf{z}}_{j}(t)$ to fluctuations in ${\bf{z}}_{k}(t)$ at Laplace mode $s$. At mode $s=0$ (the static or “DC” mode), we recover the connectivity from the steady-state or quasi-steady state analysis, $\hat{H}(0)=W$. Similarly, $\hat{B}(0)=1$. Hence, the quasi-steady state analysis from above evaluated how the DC component of the input, $\hat{{\bf{x}}}(0)$, affects the DC component of the dynamics, $\hat{{\bf{z}}}(0)$, even when we do not assume that ${\bf{x}}(t)$ changes slowly. The analysis here is a generalization of the quasi-steady state analysis to non-static modes, $s\neq 0$.
Note also that Eq. (13) from Appendix B can be derived from Eqs. (20) and (19) above by first taking $s=-2\pi if$ in Eq. (20) to switch from the Laplace domain to the Fourier domain, and then applying the Wiener–Khinchin theorem to write $\widetilde{C}^{{\bf{z}}}(f)=\widetilde{{\bf{z}}}(f)\widetilde{{\bf{z}}}^{*}(f)=\hat{{\bf{z}}}(-2\pi if)\hat{{\bf{z}}}^{*}(-2\pi if)$. Hence, the approach in this section is a generalization of the approach in Appendix B.
To interpret Eq. (20), let us consider two separate limits. At modes that are much slower than the intrinsic timescale of the dynamics, $|s|\ll 1/\tau$, we have that $\hat{H}(s)\approx W$ and $\hat{B}(s)\approx 1$, so
$\hat{{\bf{z}}}(s)\approx[I-W]^{-1}\hat{{\bf{x}}}(s)$
and we recover the results from our quasi-steady state analysis at these modes. In particular, Claim 1 holds and we expect low-rank suppression at mode $s$ whenever $W$ is strongly low rank. While this may at first seem like a repetition of our results from the quasi-steady state analysis, note that we do not need to assume that ${\bf{x}}(t)$ varies slowly here. The Laplace analysis at slow modes ($|s|\ll 1/\tau$) applies even when ${\bf{x}}(t)$ has power at slow and fast modes.
At fast modes, $|s|\gg\tau$, the identity term in Eq. (20) dominates the $\hat{H}$ and $\hat{B}$ terms so that
$\hat{{\bf{z}}}(s)\approx\hat{{\bf{x}}}(s)$
at fast modes. Therefore, fast modes (representing transient responses to transient input dynamics) do not exhibit low-rank suppression, but they do produce high-dimensional responses to high-dimensional stimuli. Modes between these two extremes interpolate between them.
In summary, when our previous assumption that ${\bf{x}}(t)$ varies more slowly than $\tau$ is violated, we expect to see low-rank suppression at slow temporal modes, but not at fast temporal modes.
Additionally, working in the Laplace domain allows us to expand our class of models. In all of the analysis above, we considered dynamics obeying the system of ordinary differential equations in Eq. (4), which we later linearized. The dynamics in Eq. (4) capture a large class of models, but are not fully general. Specifically, in Eq. (4), whenever $t_{0}<t_{1}$, the value of ${\bf{z}}(t_{1})$ is fully determined by the value of ${\bf{z}}(t_{0})$. However, some systems like integral equations, integro-differential equations, operator equations, and non-Markovian stochastic systems have a history dependence that cannot be captured by a system of ordinary differential equations like Eq. (4). In particular, we can weaken our assumptions on the dynamics of ${\bf{z}}(t)$ and assume only that ${\bf{z}}(t)$ is the response of a causal, time-invariant system that satisfies
${\bf{z}}(t)=\mathcal{F}\left(\{{\bf{z}}(s),{\bf{x}}(s)\}_{s<t}\right).$ (21)
In other words, ${\bf{z}}(t)$ is an arbitrary time-translation equivariant, causal function of the history of ${\bf{z}}(s)$ and ${\bf{x}}(s)$. Eq. (21) encompasses an extremely broad class of systems, including those described by Eq. (4).
We again begin by describing how to define a network of interactions, $W$, from the general definition of network dynamics in Eq. (21). As above, we consider a linearization around the response, ${\bf{z}}_{0}$, to a baseline input, ${\bf{x}}_{0}$. However, in this case, the baseline input and response can depend on time, ${\bf{x}}_{0}(t)$ and ${\bf{z}}_{0}(t)$. We again consider a weak perturbation away from the baseline, ${\bf{x}}_{p}(t)={\bf{x}}_{0}(t)+\epsilon{\bf{x}}(t),$ which produces a corresponding deviation
${\bf{z}}_{p}(t)=\mathcal{F}[{\bf{z}}_{p},{\bf{x}}_{p}]={\bf{z}}_{0}(t)+\epsilon{\bf{z}}(t)+\mathcal{O}(\epsilon^{2}).$
A first order Volterra expansion gives Eq. (18) as an implicit equation for ${\bf{z}}(t)$, but the matrices $\hat{H}$ and $\hat{B}$ are no longer defined as above. Instead, the matrices $\hat{H}$ and $\hat{B}$ are functional derivatives of $\mathcal{F}$, evaluated at the baseline. Under the dynamics in Eq. (6), we recover the $\hat{H}$ and $\hat{B}$ defined in Eq. (19). Hence, this approach generalizes the results in Claim 1 to fast, transient inputs and to the very general class of dynamics defined by Eq. (21).
Appendix D Additional Simulations and Data
@@
This appendix contains results from additional simulations and data. The simulations themselves and data are described in the respective captions and contextualized within the main text.
@@
Figure 11: Dynamics of a network with rank-one structure driven by an input perturbation with one dimensional structure. Same as Figure 1 except the dynamics obeyed $\tau d{\bf{z}}/dt=-{\bf{z}}+W{\bf{z}}+W_{x}\bm{\eta}(t)$ where $\bm{\eta}(t)$ is a realization of the same Gaussian stochastic process used for ${\bf{x}}(t)$ in Figure 1, and $W_{x}$ is an effectively low-rank matrix generated identically to, but independently from $W$. The “effective input”, ${\bf{x}}(t)=W_{x}\bm{\eta}(t)$, is therefore low-dimensional (see green dots in panel a). The network dynamics, ${\bf{z}}(t)$, inherit low-dimensional dynamics from ${\bf{x}}(t)$ (blue dots in panel a), but low-rank suppression and cancellation are still exhibited along the vector, $\bf{u}$}, defining the low-rank structure of $W$ (panels b–d).
Figure 12: Response properties of a network with independent singular left- and right-singular vectors. Same as Figure 2 except $W_{0}=c{\bf{u}}{\bf{v}}^{T}$ where $\bf{u}$} and $\bf{v}$} are independent random unit vectors. The network behaves similarly to the network in Figure 2 because $\bf{u}$} and $\bf{v}$} are nearly orthogonal.
@@
Figure 19: Singular values and dynamics in an epidemiological network. a) Singular values of $W$ and b) projections of $\bf{z}$} for the model in Figure 6.

### v5

@@ Appendix B Details for the linearization of dynamics around a stable equilibrium
The proof of Claim 1 also tells us the dominant directions of variability whenever dynamics are low-dimensional as in Figure 2 and Figure 3g. Since $\bf{z}$} is approximately aligned to the $U$ in the proof, we may conclude that low-dimensional dynamics are caused by excess variability along the column space of $U$, i.e., the column space of $W_{0}$. This explains the observations in Figure 2e. More specifically, low-dimensional dynamics arise when external input and the random part of internal input (i.e., $W_{1}{\bf{z}}$) have variability that overlap with the column space of $V$. Because these right singular vectors are associated with large singular values of $W$, this variability gets mapped to strong variability in the column space of $U$. Since $P$ has small singular values, this variability in the column space of $U$ does not overlap with the column space of $V$, so it cannot be cancelled by recurrent interactions. This creates an effectively feedforward mapping that amplifies variability along the directions ${\bf{v}}_{k}$ and maps it to variability in the directions of ${\bf{u}}_{k}$. This mechanism is a form of non-normal amplification [26, 27, 28] (see the discussion in Section IV.5 for a more in-depth comparison).
These results can also help us understand some subtleties of low-rank suppression. Our theory from Section IV.3 shows that the input ${\bf{x}}_{a}={\bf{u}}_{k}+(W_{1}{\bf{v}}_{k}+{\bf{v}}_{k})/\sigma_{k}$ is suppressed. Since $\sigma_{k}$ is large, this input is close to ${\bf{u}}_{k}$, but has a correction term ${\bf{x}}_{c}=(W_{1}{\bf{v}}_{k}+{\bf{v}}_{k})/\sigma_{k}$. In Figure 1g, the “aligned” input that produced a suppressed response was chosen simply as ${\bf{x}}={\bf{u}}$, without the correction term (using ${\bf{x}}_{a}$ with the correction term also produced a suppressed response; not pictured). For the example in Figure 2, we used ${\bf{x}}={\bf{x}}_{a}$ with the correction term because using ${\bf{x}}={\bf{u}}$ without a correction term did not produce a suppressed response. To understand why this happens, note that networks like the one in Figure 2 (for which $P=V^{T}U$ is singular or has a small singular value) take inputs from the directions ${\bf{v}}_{k}$ and amplify them in proportion to $\sigma_{k}$ (according to our proof of Claim 1). Hence, the ${\bf{v}}_{k}/\sigma_{k}$ term in ${\bf{x}}_{a}$ is small by a factor of $\sigma_{k}$, but the response is also amplified by a factor of $\sigma_{k}$, so the correction term produces a $\mathcal{O}(1)$ response and cannot be ignored. In summary, the correction term, ${\bf{x}}_{c}$, is necessary to include when $P$ is singular or has small singular values (as in Figure 2).
Appendix E Dimensionality of dynamics when inputs are not slowly varying.
In Section IV.5 and Appendix D, we assumed that ${\bf{x}}(t)$ varied on a slower timescale than the timescale, $\tau$, of the network dynamics, justifying the quasi-steady state approximation in Eq. (6). While this assumption simplifies our analysis and conclusions, we can understand the dimensionality of dynamics without it. Here, we analyze the dimensionality of ${\bf{z}}(t)$ under all of the conditions from Assumptions 1–3 except the assumption that ${\bf{x}}(t)$ is slowly varying.
Beginning with Eq. (14), we see that the dimensionality of ${\bf{z}}(t)$ in general depends on the interaction between $\widetilde{C}^{{\bf{x}}}(f)$, $W$ and $r(f)$ across different frequencies, $f$. Here, we again focus on the special case that each ${\bf{x}}_{j}(t)$ is i.i.d. In this case, $\widetilde{C}^{{\bf{x}}}(f)=c_{x}(f)I$ where $c_{x}(f)$ is the scalar power spectral density of each ${\bf{x}}_{j}(t)$. Therefore, the covariance matrix of ${\bf{z}}(t)$ can be written as [42]
$\overline{C}^{{\bf{z}}}=\int_{-\infty}^{\infty}[r(f)I-W]^{-1}[r(f)I-W]^{-*}c_{x}(f)df$ (18)
where $\cdot^{-*}$ is the inverse conjugate transpose and $r(f)=1-2\pi\tau fi$.
First consider $\widetilde{C}^{{\bf{z}}}(f)$ at frequencies lower than the timescale of the dynamics, $f\ll 1/(2\pi\tau)$. At these frequencies, we have that $r(f)\approx 1$ so that
$\widetilde{C}^{{\bf{z}}}(f)\approx[I-W]^{-1}[I-W]^{-T}c_{x}(f)$
at sufficiently low frequencies. Since $c_{x}(f)$ is a scalar, we have that
$\overline{C}^{{\bf{z}}}\propto[I-W]^{-1}[I-W]^{-T}$
for $f\ll 1/(2\pi\tau)$. Therefore, the arguments and conclusions from Section IV.5, Appendix D, and Claim 1 apply in general to the dimensionality of dynamics at low frequencies. Specifically, the dimensionality of the dynamics at low frequencies (i.e., the slow dynamics) can only be low-dimensional when $P=V^{T}U$ has a small singular value. This is true even when ${\bf{x}}(t)$ has power at high frequencies.
However, the stationary covariance matrix of ${\bf{z}}(t)$ accounts for variability at all frequencies (see Eqs. (18)). When ${\bf{x}}(t)$ is slowly varying, $c_{x}(f)\approx 0$ at higher frequencies (see Figure 13), so the low-frequency analysis above is sufficient (consistent with the results obtained in Appendix D and Section IV.5). When ${\bf{x}}(t)$ has power at high frequencies, these higher frequencies need to be accounted for.
At high frequencies satisfying $f\gg\|W\|/(2\pi\tau)$, the $r(f)I$ terms in Eq. (18) dominates the $W$ terms so that
$\widetilde{C}^{{\bf{z}}}(f)\approx\frac{c_{x}(f)}{1+4\pi^{2}f^{2}\tau^{2}}I$
which is a multiple of the identity and therefore full-rank. Hence, when ${\bf{x}}(t)$ has variability at these high frequencies (i.e., when $c_{x}(f)$ is not close to zero for $f\gg\|W\|/(2\pi\tau)$; see Figure 14a), these high frequency components can only add high-dimensional variability to ${\bf{z}}(t)$ because ${\bf{z}}(t)$ tracks ${\bf{x}}(t)$ at high frequencies.
If ${\bf{x}}(t)$ has variability only at these high frequencies ($f\gg\|W\|/(2\pi\tau)$), then only these frequencies would contribute to the integral in Eq. (18) and therefore ${\bf{z}}(t)$ would be high-dimensional regardless of the structure of $W$. However, it is somewhat rare that natural stochastic processes have power only at higher frequencies.
More commonly, when ${\bf{x}}(t)$ varies quickly, it has power at both low and high frequencies. In other words, $c_{x}(f)$ is far from zero for both small and large $f$. For example, white noise has equal power at all frequencies and Orenstein Uhlenbeck processes have more power at low than high frequencies even when they have fast timescales. When ${\bf{x}}(t)$ has power at low and high frequencies, the integral in Eq. (18) has contributions from both low and high frequencies. Low-dimensional variability can only be introduced at low frequencies and, by Claim 1 and our arguments above, can only be introduced when $P=V^{T}U$ has a small singular value. Since these low-dimensional dynamics represent an asymptotically large singular value, this combination of low-dimensional variability at low frequencies and high-dimensional variability at high frequencies results in low-dimensional variability overall since the low-dimensional variability is strong while the high-dimensional variability is $\mathcal{O}(1)$. As a result, when ${\bf{x}}(t)$ has power at both low and high frequencies, the conclusions of Claim 1 apply. Specifically, low-dimensional variability arises only when $P$ has small singular values. To test this conclusion, we repeated the simulations from Figure 3 with timescales swapped. Specifically, we set $\tau=10$ and $\tau_{x}=1$ so that $c_{x}(f)$ is far from zero at both low frequencies ($f\ll 1/(2\pi\tau)$) and high frequencies ($f\gg\|W\|/(2\pi\tau)$; see Figure 14a in Appendix F). As predicted, low-dimensional variability only arises in the case that $P=V^{T}U$ has a singular value at zero (Figure 14b–f), identical to the case when ${\bf{x}}(t)$ varies slowly (compare Figure 14 to Figure 3).
Appendix F Additional Simulations and Data
@@
This appendix contains results from additional simulations and data. The simulations themselves and data are described in the respective captions and contextualized within the main text and Appendix.
@@
Figure 8: Dynamics of a network with rank-one structure driven by an input perturbation with one dimensional structure. Same as Figure 1 except the dynamics obeyed $\tau d{\bf{z}}/dt=-{\bf{z}}+W{\bf{z}}+W_{x}\boldsymbol{\eta}(t)$ where $\boldsymbol{\eta}(t)$ is a realization of the same Gaussian stochastic process used for ${\bf{x}}(t)$ in Figure 1, and $W_{x}$ is an effectively low-rank matrix generated identically to, but independently from $W$. The “effective input”, ${\bf{x}}(t)=W_{x}\boldsymbol{\eta}(t)$, is therefore low-dimensional (see green dots in panel a). The network dynamics, ${\bf{z}}(t)$, inherit low-dimensional dynamics from ${\bf{x}}(t)$ (blue dots in panel a), but low-rank suppression and cancellation are still exhibited along the vector, $\bf{u}$}, defining the low-rank structure of $W$ (panels b–d).
Figure 9: Response properties of a network with independent left- and right-singular vectors. Same as Figure 2 except $W_{0}=c{\bf{u}}{\bf{v}}^{T}$ where $\bf{u}$} and $\bf{v}$} are independent random unit vectors. The network behaves similarly to the network in Figure 2 because $\bf{u}$} and $\bf{v}$} are nearly orthogonal.
@@
Figure 12: Dynamics of a stable network with rank-one structure close to criticality. Same as Figure 1 except $\rho=0.95$. The fixed point at ${\bf{z}}=0$ is stable, but close to instability because the maximum real part of the eigenvalues of the Jacobian is close to 0.
Figure 13: Power spectral density of ${\bf{x}}(t)$ used in simulations. Average power spectral density of ${\bf{x}}_{j}(t)$ (green) compared to $|r(f)|^{2}=1/(1+4\pi^{2}f^{2}\tau^{2})$ (gray) when $\tau_{x}=10$ and $\tau=1$, as in all figures except for Figure 14. Vertical red line shows $f=1/(2\pi\tau)$.
Figure 14: High- and low-dimensional dynamics when inputs are not slow. Simulations from Figure 3 were repeated with $\tau_{x}=1$ and $\tau=10$. a) Average power spectral density of ${\bf{x}}_{j}(t)$ (green) compared to $|r(f)|^{2}=1/(1+4\pi^{2}f^{2}\tau^{2})$ (gray). Vertical red line shows $f=1/(2\pi\tau)$. Since ${\bf{x}}(t)$ is faster than intrinsic network dynamics ($\tau_{x}<\tau$), $c_{x}(f)$ has power at higher frequencies where $r(f)\approx 0$. b-e) Same as Figure 3 panels c,f,i,l,o respectively except that we set $\tau_{x}=1$ and $\tau=10$. Variability in ${\bf{z}}(t)$ is high- and low-dimensional under the same conditions.
@@
Figure 19: Singular values and dynamics in an epidemiological network. a) Singular values of $Q$ (red) and $W$ (blue). b) Projections of $\bf{z}$} for the model in Figure 6. c) Same as Figure 6a except the dominant singular vector of $Q$ was used to define the aligned direction instead of the dominant singular value of $W$. The lack of suppression demonstrates the necessity of using the effective connectivity matrix, $W$, in place of $Q$, which is the connectivity matrix used to define the dynamics.
Figure 20: Response properties of a weakly low-rank network with $\rho\ll 1$ and $c<0$. Same as Figure 1 except c=0.5c=-0.5 and $\rho=0.01$.
Figure 21: Response properties of a weakly low-rank network with $\rho\ll 1$ and $c>0$. Same as Figure 1 except $c=0.5$ and $\rho=0.01$. Note that stability requires that $c<1$ when $c>0$.
Figure 22: Quantifying low-rank suppression as the maximum singular value is scaled. We repeated the simulation from Figure 1 while changing the value of $c<0$ to change the value of the maximum singular value of $W$, $\sigma_{\textrm{max}}=|c|$. This demonstrates a smooth transition between weakly and strongly low-rank structure. For each value of $c$, we computed the ratio of $\|{\bf{z}}\|$ in response to an aligned input versus a random input, i.e., the height of the two bumps in Figure 1g as a measure of low-rank suppression. Low-rank suppression occurs when this ratio is small.
Figure 23: High-dimensional dynamics in weakly low-rank networks. a) Singular values for a weakly low-rank network in which $\rho=0.01$, $\sigma_{1}=0.5$, and $\sigma_{2}=0.25$. b-f) Same as Figure 3c,f,i,l,o except that $\rho=0.01$, , $\sigma_{1}=0.5$, and $\sigma_{2}=0.25$. The network response is high-dimensional in every case.